OpenStudy (anonymous):

Which value for x makes the open sentence true? ill give medal and fan plz need help!!!!! 6 • 7 – 5x + x2 ≥ x3 + x2

OpenStudy (anonymous):

Is this the inequality you refer to?\[6 \times 7 - 5x + 2x \geq 3x + 2x\]

Mehek (mehek14):

what are the values that you have?

OpenStudy (anonymous):

forgot to write dow the choices here u go 3 4 6 5

OpenStudy (anonymous):

3 4 6 and 5

OpenStudy (anonymous):

plz help

Mehek (mehek14):

start by plugging in the vales

OpenStudy (anonymous):

iv tried and nothing came out right thats why i cam on to open study

Mehek (mehek14):

ok it's not 3

Mehek (mehek14):

not 4

OpenStudy (anonymous):

the closet i got was either 4 or 5

Mehek (mehek14):

not 6

Mehek (mehek14):

I think you should double check with your teacher

Mehek (mehek14):

because none are true

OpenStudy (anonymous):

which one would be closet

OpenStudy (anonymous):

The simplified constraint on x is given by \[42 \geq 8x, \quad \text{or } x \leq 42 / 8\]

OpenStudy (anonymous):

And 42/8 is 21/4, which is 5 and 1/4

OpenStudy (anonymous):

So 3, 4, and 5 satisfy this inequality, with 5 being closest to the bound 5 and 1/4

OpenStudy (anonymous):

the answer was six i missed a step but thank anyways guys for all the help